An aqueous solution containing iron (II) (Fe2+) at a concentration of 2.5 x 10-3 mg cm-3 is subjected to the above treatment and measures an absorbance of 0.524. A water sample of unknown iron (II) (Fe2+) concentration is similarly treated and measures an absorbance of 0.350. Calculate the concentration of iron (II) (Fe2+) in the unknown sample in mg cm-3.


1
Expert's answer
2021-03-23T04:03:46-0400

According to the equation:

A = ecl

where e - molar absorption coefficient, c - concentration, l - optical path length, A - absorbance.

As molar absorption coefficient and optical path length are the same for both solutions, the concentration equals:

C2 = A2 / el = A2 × C1 / A1 = 0.350 × 2.5 x 10-3 mg cm-3 / 0.524 = 1.7 × 10-3 mg cm-3


Answer: 1.7 × 10-3 mg cm-3